Assume the demand function for Dodo Corporation is expressed as:

Q x = 500 – 2P x + 10P c + 5Y
Q x = quantity demanded in thousand
P x = price
P c = price of a related good C
Y = income per capita in thousand

a. If P c = RM2 and Y = RM10, derive the demand curve as price function of quantity.
b. Is good X a normal or inferior good? Why?
c. What would you expect to happen on the sales of this company if the competitor launched a
price cutting policy? Why?
Answer #1

x = 50 2 2 + lopet sy put Pc = 2 and y z lo - lx - 5w - 2Px + 10 (2) +5 (10) - Qx = sw-2Px + 20+50 10x = 570-2PX (Demond fon